### v1

# Testing by Betting while Borrowing and Bargaining

Abstract
Testing by betting has been a cornerstone of the game-theoretic statistics literature. In this framework, a betting score (or more generally an e-process), as opposed to a traditional p-value, is used to quantify the evidence against a null hypothesis: the higher the betting score, the more money one has made betting against the null, and thus the larger the evidence that the null is false. A key ingredient assumed throughout past works is that one cannot bet more money than they currently have. In this paper, we ask what happens if the bettor is allowed to borrow money after going bankrupt, allowing further financial flexibility in this game of hypothesis testing. We propose various definitions of (adjusted) evidence relative to the wealth borrowed, indebted, and accumulated. We also ask what happens if the bettor can “bargain”, in order to obtain odds bettor than specified by the null hypothesis. The adjustment of wealth in order to serve as evidence appeals to the characterization of arbitrage, interest rates, and numéraire-adjusted pricing in this setting.
1 Introduction
Aggressive gamblers put borrowed money on the table to increase their potential win. Risk-seeking traders enter leveraged positions for exposure to higher expected returns. In this paper, we study analogous situations in game-theoretic statistics.
In the standard setup of game-theoretic statistics, bets are offered to a statistician to test the verity of a null hypothesis [Shafer, 2021]. Before each outcome is revealed, the statistician places an arbitrary fraction of their wealth to bet on its value, at odds specified by the null hypothesis. The total wealth of the gambler can be interpreted as the amount of evidence available against the null.
In particular, we consider the mean estimation set-up studied by various authors including Shafer and Vovk [2005, Section 3] and Waudby-Smith and Ramdas [2024], where there is an infinite sequence of bounded random variables with a common conditional mean $\mu$. We are chiefly concerned with the special case where these random variables are coin tosses; that is, they are i.i.d. drawn from some distribution on {±$\{\pm 1\}$. The fairness of the coin, $H_{0}:\mu=1/2$, is to be tested. Extending our exposition to the general bounded case with a common conditional mean is straightforward.
It is worth noting that an overarching assumption has been made by the entirety of the previous game-theoretic probability literature. Namely, the bettor is not allowed to risk in the game more than what they have. For example, in the prototypical game of “fair coin” (which we model our set-up after), Shafer and Vovk [2005, Section 3.1] allow the bettor to freely choose the volume of bets but they lose the game immediately if their wealth becomes negative, essentially banning any bets larger than the current wealth. The same “losing at bankruptcy” rule applies to the entire book by these authors, as well as in their follow-up book [Shafer and Vovk, 2019]. This constraint has essentially remain unchanged and unchallenged in the game-theoretic statistics literature [Ramdas et al., 2023].
In this paper, however, we allow the betting statistician to borrow interest-free before placing each bet without being forced to quit the game in the possible event of indebtedness, and study if the emergence of evidence is changed in any way correspondingly. Further, we discuss the related scenario where the bets are underpriced such that arbitrage is possible without evidence.
Throughout the paper, we consider a fixed measurable space $(\Omega,\mathcal{A})$ with measurable functions $X_{1},X_{2},\dots:\Omega\to\{\pm 1\}$, construed as repeated coin tosses; and denote by $\{\mathcal{F}_{t}\}$ the canonical filtration they generate, i.e., $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$. We denote by $\mathbb{P}_{p}$ the probability measure on $(\Omega,\mathcal{A})$ such that $X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}\operatorname{Rad}_{p}$ (Rademacher ±$\pm 1$-valued random variables with probability of 1 being $p$), and $\mathbb{E}_{p}$ the corresponding expectation or conditional expectation.
2 Review of Testing by Betting (without Borrowing)
The classical set-up of testing by betting pioneered by Shafer and Vovk [2005] involves the following game with two players, Casino [^1] This player is referred to as Forecaster in the original work, but in our work, Casino seems a better analogy. and Statistician. Casino tosses a coin repeatedly, realized as the random variables $X_{1},X_{2}\dots$ mentioned above, and Statistician does not believe that the coin is fair. At each time $t=1,2,\dots$, Casino offers a double-or-nothing bet to Statistician on the upcoming coin toss $X_{t}$. Statistician is allowed to place any fraction (between 0 and 1) of their wealth on either heads or tails. Formally speaking, let $\{\lambda_{t}\}_{t\geqslant 1}$ be a process bounded in $[-1,1]$, predictable with respect to the filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$, where $\lambda_{t}$ the fraction of wealth Statistician bets on $X_{t}=1$. Let $W_{0}=1$ be the initial wealth of Statistician, whose wealth process therefore evolves as
$W_{t}=W_{t-1}\cdot(1+\lambda_{t}X_{t})$ (1)
over the course of the game.
Under $\mathbb{P}_{1/2}$, i.e., when the coin is fair, $\mathbb{E}_{1/2}(X_{t})=0$, implying that the wealth process $\{W_{t}\}$ is a nonnegative martingale. Ville’s inequality [Ville, 1939] then implies that, for all $x>1$,
$\mathbb{P}_{1/2}\left(\sup_{t}W_{t}\geqslant x\right)\leqslant 1/x;$ (2)
or equivalently, for any $\{\mathcal{F}_{t}\}$-stopping time $\tau$ (possibly infinite),
$\mathbb{P}_{1/2}\left(W_{\tau}\geqslant x\right)\leqslant 1/x.$ (3)
Along with the optional stopping theorem for nonnegative martingales, which states that $\mathbb{E}_{1/2}W_{\tau}\leqslant 1$ for any stopping time $\tau$, both inequalities above lend credibility to the interpretation that Statistician’s wealth $W_{t}$ can be seen as the evidence against the fairness of the coin $\mathbb{P}_{1/2}$. Inequality (2) states that if the coin was fair, it would be unlikely for Statistician ever to be wealthy, and thus a large value of $W_{t}$ at any time means that the coin is unlikely to be fair. Inequality (3) states that Statistician may employ any stopping rule and a large final stopped wealth discredits $\mathbb{P}_{1/2}$. Related to both is the concept of e-values, which are nonnegative random variables with expected values at most 1. In particular, $W_{\tau}$ is an e-value for any stopping time $\tau$.
The scenario described in this section is a classical motivating example of game-theoretic statistics, where traditional notions in statistics like p-values are eschewed in favor of e-values like $W_{\tau}$. In what follows, we begin our attempt to generalize these concepts to the case where borrowing is allowed.
3 Evidence in Borrowed Betting
We now suppose before placing the bet on $X_{t}$ at time $t$, Statistician borrows [^2] It does not matter whom Statistician borrows the money from. $\beta_{t}$ amount of money, increasing (or decreasing, if $\beta_{t}<0$) their current (gross) wealth $W_{t-1}$ to $W_{t-1}+\beta_{t}$, of which Statistician then bets a $\lambda_{t}$ fraction on $X_{t}=1$. Formally, let $\{\beta_{t}\}_{t\geqslant 1}$ and $\{\lambda_{t}\}_{t\geqslant 1}$ be two predictable processes on $\{\mathcal{F}_{t}\}_{t\geqslant 0}$. We assume $\lambda_{t}\in[-1,1]$ and $\beta_{t}\geqslant-W_{t-1}$ for each $t$. In particular, we allow $\beta_{t}$ to be negative which could correspond to Statistician “paying back” earlier debts. Now, Statistician’s wealth process is defined as
$W_{0}=1,\quad W_{t}=(W_{t-1}+\beta_{t})\cdot(1+\lambda_{t}X_{t}),$ (4)
a nonnegative process adapted to $\{\mathcal{F}_{t}\}$.
To further clarify the set-up, we introduce the liabilities process denoting the total borrowed amount up to time $t$,
$L_{t}=\sum_{i=1}^{t}\beta_{i};$ (5)
as well as the net wealth process for the difference between “gross” wealth and liabilities,
$N_{t}=W_{t}-L_{t}.$ (6)
Recall that in the previous setting without borrowing in Section 2, the wealth process is a nonnegative martingale under $\mathbb{P}_{1/2}$. In our current case with borrowing, we have the following proposition on the more general relation between these three processes.
Proposition 3.1 (Doob decomposition of wealth).
The net wealth process $\{N_{t}\}$ is a martingale on $\{\mathcal{F}_{t}\}$ under $\mathbb{P}_{1/2}$. Further, $W_{t}=N_{t}+L_{t}$ is the Doob decomposition of the process $\{W_{t}\}$ into the martingale $\{N_{t}\}$ and the predictable process $\{L_{t}\}$. Consequently, $\{W_{t}\}$ is a submartingale if and only if the amount borrowed each time $\beta_{t}\geqslant 0$.
Proof.
$\mathbb{E}_{1/2}(N_{t}|\mathcal{F}_{t-1})=(W_{t-1}+\beta_{t})\cdot\mathbb{E}_{1/2}(1+\lambda_{t}X_{t}|\mathcal{F}_{t-1})-L_{t}=W_{t-1}+\beta_{t}-L_{t}=N_{t-1}$, therefore $\{N_{t}\}$ is a $\mathbb{P}_{1/2}$-martingale. It is easy to see that $\{L_{t}\}$ is predictable and $L_{0}=0$. Therefore, $W_{t}=N_{t}+L_{t}$ is the Doob decomposition of $\{W_{t}\}.$ The final claim is immediate. ∎
We would like to study how Statistician, now holding both wealth $W_{t}$ and liabilities $L_{t}$, can quantify the evidence against $\mathbb{P}_{1/2}$, based on either $W_{t}$ or $N_{t}$ or both. To generalize our previous discussion on interpreting evidence via inequalities (2) and (3), let us introduce the following definitions.
Definition 3.2 (Tail evidence).
Let $\tau$ be a stopping time on $\{\mathcal{F}_{t}\}$ and $a>0,b\geqslant 0$ be constants. We say that an $\mathcal{F}_{\tau}$-measurable random variable $E$ is an $(a,b)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$ if the tail bound
$\mathbb{P}_{1/2}(E\geqslant x)\leqslant\frac{a}{x-b}$ (7)
holds for all $x>b$.
Random variables that are measurable with respect to the stopped $\sigma$-algebra $\mathcal{F}_{\tau}$ are often constructed by stopping at $\tau$ a process adapted to $\{\mathcal{F}_{t}\}$. This definition of tail evidence above follows the same reasoning that “if $\mathbb{P}_{1/2}$ holds, $E$ is unlikely to be large” as discussed in Section 2; the e-value $W_{\tau}$ back in (3), per Definition 3.2, is a $(1,0)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$, and the constants $a,b$ in the definition are seen as “discount factors” of the evidence value $E$. More generally, if $a^{-1}(E-b)$ is an e-value, then $E$ is an $(a,b)$-tail evidence. Conversion from an $(a,b)$-tail evidence to a p-value for testing the null $\mathbb{P}_{1/2}$ can be done via $E\mapsto\frac{a}{(E-b)^{+}}$.
We further define the following sequential, stopping time-free concept of evidence via a time-maximal inequality like (2).
Definition 3.3 (Sequential tail evidence).
If a process $\{E_{t}\}$ adapted to $\{\mathcal{F}_{t}\}$ satisfies, for all $x>b$,
$\mathbb{P}_{1/2}\left(\sup_{t}E_{t}\geqslant x\right)\leqslant\frac{a}{x-b},$ (8)
we say that it is a $(a,b)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
For example, Ville’s inequality states that any nonnegative $\mathbb{P}_{1/2}$-supermartingale $\{M_{t}\}$ is an $(M_{0},0)$-sequential tail evidence. The stopping time-dependent Definition 3.2 can be converted to the time-maximal Definition 3.3, if the same constant pair $(a,b)$ works for all bounded stopping times, as shown in the following proposition.
Proposition 3.4.
Let $\{E_{t}\}$ be a process adapted to $\{\mathcal{F}_{t}\}$. It is an $(a,b)$-sequential tail evidence against $\mathbb{P}_{1/2}$ if and only if for any bounded stopping time $\tau$, $E_{\tau}$ is an $(a,b)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$.
Proof.
First, suppose $E_{\tau}$ is an $(a,b)$-tail evidence at $\tau$ for any bounded stopping time $\tau$. Let $\sigma$ be the stopping time $\inf\{t:E_{t}\geqslant x\}$, where by convention $\inf\varnothing=\infty$. Then, since $E_{\sigma\wedge T}$ is an $(a,b)$-tail evidence against $\mathbb{P}_{1/2}$ at $\sigma\wedge T$,
$\mathbb{P}_{1/2}\left(\sup_{t\leqslant T}E_{t}\geqslant x\right)=\mathbb{P}_{1/2}(E_{\sigma\wedge T}\geqslant x)\leqslant\frac{a}{x-b}.$ (9)
Since the same probability bound holds for upward nested events $\{\sup_{t\leqslant T}E_{t}\geqslant x\}$ with $T=1,2,\dots$, we see that
$\mathbb{P}_{1/2}\left(\sup_{t}E_{t}\geqslant x\right)=\mathbb{P}_{1/2}\left(\bigcup_{T=1}^{\infty}\left\{\sup_{t\leqslant T}E_{t}\geqslant x\right\}\right)\leqslant\frac{a}{x-b},$ (10)
concluding that $\{E_{t}\}$ is an $(a,b)$-sequential tail evidence.
Next, suppose $\{E_{t}\}$ is an $(a,b)$-sequential tail evidence and let $\tau\leqslant T$ be a bounded stopping time. Then $\mathbb{P}_{1/2}(E_{\tau}\geqslant x)\leqslant\mathbb{P}_{1/2}(\sup_{t\leqslant T}E_{t}\geqslant x)\leqslant\frac{a}{x-b}$, concluding that $E_{\tau}$ is an $(a,b)$-tail evidence at $\tau$. ∎
The proof of Proposition 3.4 above, we note, is itself a proof of Ville’s inequality by setting $b=0$, combined with the optional stopping theorem. We refer the reader to Howard et al. [2021, Lemma 3] for a stronger result on the equivalence of stopped probability and time-maximal probability bounds.
In the rest of the paper, we shall present several evidences and sequential evidences based on Statistician’s wealth $W_{t}$ or net wealth $N_{t}$. As we shall see soon, these are possible under any of the several additional assumptions we shall make on the amount borrowed each time $\beta_{t}$. The following are some general assumptions that may (or may not) be invoked. First, we may assume that Statistician cannot pay back their debts.
Assumption 3.5 (Positivity of borrowings).
For every $t$, $\beta_{t}\geqslant 0$.
Second, to find evidences at a stopping time $\tau$, we may need either $\tau$ or the net wealth process to be “well-behaved”.
Assumption 3.6.
The net wealth martingale $\{N_{t}\}$ satisfies the supermartingale optional stopping theorem with respect to stopping times $\tau$ and 0; that is, $\mathbb{E}_{1/2}N_{\tau}\leqslant N_{0}=1$.
For example, Assumption 3.6 is satisfied if any of the following conditions hold: (1) $\tau$ is bounded; (2) $\{N_{t}\}$ is uniformly integrable and $\tau$ is finite; (3) $\{N_{t}\}$ is bounded from below.
4 Gross Wealth as Evidence
We first establish conditions on which Statistician’s gross wealth $\{W_{t}\}$, with liabilities undeducted, can act as evidence against $\mathbb{P}_{1/2}$.
The following assumptions restrict Statistician to borrowing only up to, in expectation, a fixed amount of money, either at some stopping time or at any fixed time.
Assumption 4.1 (Bounded expected stopped liabilities).
The random variable $L_{\tau}$ is integrable under $\mathbb{P}_{1/2}$: $L=\mathbb{E}_{1/2}L_{\tau}<\infty$.
Assumption 4.2 (Bounded expected liabilities).
The process $\{L_{t}\}$ is bounded in expectation under $\mathbb{P}_{1/2}$: $L=\sup_{t}\mathbb{E}_{1/2}L_{t}<\infty$.
Then, we have the following results.
Proposition 4.3.
(i) Suppose a stopping time $\tau$ satisfies both Assumptions 3.6 and 4.1. Then, $(1+L)^{-1}W_{\tau}$ is an e-value under $\mathbb{P}_{1/2}$, thus the wealth $W_{\tau}$ is a $(1+L,0)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$. (ii) Under Assumptions 3.5 and 4.2, the wealth process $\{W_{t}\}$ is a $(1+L,0)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
Proof.
(i) Note that $W_{\tau}$ is a nonnegative random variable and $\mathbb{E}_{1/2}W_{\tau}=\mathbb{E}_{1/2}N_{\tau}+\mathbb{E}_{1/2}L_{\tau}\leqslant 1+L$. Therefore, by Markov’s inequality, $\mathbb{P}_{1/2}(W_{\tau}\geqslant x)\leqslant x^{-1}(1+L)$.
(ii) Since Assumption 3.5 holds, we learn from Proposition 3.1 that the wealth process $\{W_{t}\}$ is a nonnegative submartingale. Therefore, letting $\tau\leqslant T$ be any bounded stopping time, the optional stopping theorem implies that $\mathbb{E}_{1/2}W_{\tau}\leqslant\mathbb{E}_{1/2}W_{T}=\mathbb{E}_{1/2}N_{T}+\mathbb{E}_{1/2}L_{T}\leqslant N_{0}+L=1+L$. Therefore by Markov’s inequality, $\mathbb{P}_{1/2}(W_{\tau}\geqslant x)\leqslant x^{-1}(1+L)$, concluding the proof. ∎
Proposition 4.3 states that the total gross wealth that Statistician gathers over the bets counts as evidence against $\mathbb{P}_{1/2}$, discounted by a factor that corresponds to the average liabilities over all possible coin-tosses. It is easy to see that we can slightly weaken the assumptions of Part (ii) of Proposition 4.3, allowing Statistician to pay back debts but placing instead an upper bound on the expected total actual borrowings.
Proposition 4.4.
If
$B=\sup_{t}\mathbb{E}_{1/2}\sum_{i=1}^{t}\beta_{i}^{+}<\infty,$ (11)
the wealth process $\{W_{t}\}$ is a $(1+B,0)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
Proof.
Consider the increasing process $B_{t}=\sum_{i=1}^{t}\beta_{i}^{+}$ and $V_{t}=N_{t}+B_{t}$. Then, $V_{t}\geqslant W_{t}$ and $\{V_{t}\}$ is a submartingale. Letting $\tau\leqslant T$ be any bounded stopping time, the optional stopping theorem implies that $\mathbb{E}_{1/2}W_{\tau}\leqslant\mathbb{E}_{1/2}V_{\tau}\leqslant\mathbb{E}_{1/2}V_{T}=\mathbb{E}_{1/2}N_{T}+\mathbb{E}_{1/2}B_{T}\leqslant 1+B$. Markov’s inequality then completes the proof. ∎
There, we note, lies the key drawback of both statements in Proposition 4.3, as well as the generalized Proposition 4.4, that the discount factor of the evidence $L$ (or $B$) depends on the unobserved and unrealized counterfactual values of Statistician’s liabilities. With Part (i) of Proposition 4.3, for example, Statistician has borrowed $L_{\tau(\omega)}(\omega)$ at time $\tau$; the values $L_{\tau(\omega^{\prime})}(\omega^{\prime})$ for $\omega^{\prime}\neq\omega$, i.e., what Statistician could have borrowed in alternate worlds, somehow still discount Statistician’s evidence in the current world. Part (ii) of Proposition 4.3 has the more unpalatable feature of penalizing the evidence based on what Statistician will borrow in the future, current and alternate worlds alike. A solution is to be offered in the next section.
We finally remark that the wealth process $\{W_{t}\}$ in Proposition 4.4 is an instance of “nonnegative almost supermartingales” studied by Robbins and Siegmund [1971], and Proposition 4.4 can also be derived by letting the upper time horizon $m\to\infty$ in the maximal inequality of Robbins and Siegmund [1971, Proposition 2], which we shall quote as Proposition 8.2 in Section 8 later when discussing another extension of our results.
5 Net Wealth as Evidence
Seeing the drawbacks of using (adjusted) gross wealth as evidence, we now ask the natural question of whether the net wealth of Statistician, $\{N_{t}\}$, can be a better measure of evidence when borrowing is allowed. We now assume that the net wealth, either stopped or uniformly, is lower bounded by a constant.
Assumption 5.1 (Stopped net wealth bounded from below).
The random variable $N_{\tau}$ and non-random number $N_{\min}<1$ satisfy $N_{\tau}\geqslant N_{\min}$.
Assumption 5.2 (Net wealth bounded from below).
There exists a non-random number $N_{\min}<1$ such that $N_{t}\geqslant N_{\min}$ for all $t$.
These assumptions, we note, are easier to verify or satisfy compared to Assumptions 4.1 and 4.2. For example, if a net wealth lower bound $N_{\min}$ is set in advance, Statistician can simply restrict the amount they borrow and bet according to their current wealth and liabilities,
$(W_{t-1}+\beta_{t})(1-\lambda_{t})-\beta_{t}\geqslant L_{t-1}+N_{\min},$ (12)
to meet either assumption. These assumptions, we note, are also inherently more liberal compared to Assumptions 4.1 and 4.2, since the borrowable amount (to meet the assumption) now grows as Statistician’s current net wealth does, instead of being constrained by a fixed line of credit. We have the following results on net wealth as evidence.
Proposition 5.3.
(i) Suppose a stopping time $\tau$ satisfies both Assumptions 3.6 and 5.1. Then, $(N_{\tau}-N_{\min})/(1-N_{\min})$ is an e-value under $\mathbb{P}_{1/2}$, thus the net wealth $N_{\tau}$ is a $(1-N_{\min},N_{\min})$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$. (ii) Under Assumption 5.2, the net wealth process $\{N_{t}\}$ is a $(1-N_{\min},N_{\min})$-sequential tail evidence against $\mathbb{P}_{1/2}$.
Proof.
(i) follows from $\mathbb{E}_{1/2}(N_{\tau}-N_{\min})=1-N_{\min}$, and $N_{\tau}-N_{\min}\geqslant 0$; and (ii) follows from the fact that $\{N_{t}-N_{\min}\}$ is a nonnegative martingale. ∎
Comparing Proposition 5.3 to Proposition 4.4, we see that both results allow $\beta_{t}$ to be negative. In Proposition 4.4, there appears no benefit for Statistician to pay back debts; they may as well not pay them back by setting $\beta_{t}=0$ as it does not affect whatsoever the evidence and its discount factor $B$, which is defined only through the positive part $\beta_{t}^{+}$. In Proposition 5.3, however, evidence does benefit from paying back debts in some cases, as it increases $N_{\min}$ all other actions being equal.
A more flexible bound can be obtained by considering an “imbalanced” mixture of wealth and liabilities with the latter receiving more weight, which we call a sub-net wealth. Formally, we consider a predictable process $\eta_{t}\geqslant 1$ and the corresponding sub-liabilities process $\tilde{L}_{t}=\sum_{i=1}^{t}\eta_{i}\beta_{i}$. The sub-net wealth process is now defined as $\tilde{N}_{t}=W_{t}-\tilde{L}_{t}$, leading to the following refinements of Assumptions 5.1 and 5.2.
Assumption 5.4 (Stopped sub-net wealth bounded from below).
The random variable $\tilde{N}_{\tau}$ and non-random number $G<1$ satisfy $N_{\tau}\geqslant G$.
Assumption 5.5 (Sub-net wealth bounded from below).
There exists a non-random number $G<1$ such that $\tilde{N}_{t}\geqslant G$ for all $t$.
We have the following generalization of Proposition 5.3.
Proposition 5.6.
(i) Suppose a bounded stopping time $\tau$ satisfies Assumption 5.4. Then, $(\tilde{N}_{\tau}-G)/(1-G)$ is an e-value under $\mathbb{P}_{1/2}$, thus the sub-net wealth $\tilde{N}_{\tau}$ is a $(1-G,G)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$. (ii) Under Assumption 5.5, the sub-net wealth process $\{\tilde{N}_{t}\}$ is a $(1-G,G)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
Proof.
It suffices to show that the sub-net wealth process $\{\tilde{W}_{t}\}$ is a supermartingale. Note that $L_{t}-L_{t-1}=\beta_{t}\leqslant\eta_{t}\beta_{t}=\tilde{L}_{t}-\tilde{L}_{t-1}$, so
$\mathbb{E}_{1/2}(W_{t}-\tilde{L}_{t}|\mathcal{F}_{t-1})=\mathbb{E}_{1/2}(W_{t}-L_{t}|\mathcal{F}_{t-1})+(L_{t}-\tilde{L}_{t})=(W_{t-1}-L_{t-1})+(L_{t}-\tilde{L}_{t})$ (13)
$\leqslant$ $W_{t}-\tilde{L}_{t-1},$ (14)
concluding the proof. ∎
Intuitively speaking, Proposition 5.6 states that if Statistician is able to control the loss in a way that is stricter than lower bounding the net wealth, a possibly stronger measure of evidence is at hand. It might seem surprising that defining a process strictly less than the net wealth leads to possibly stronger evidence. We illustrate this by the following simple example and conduct a fair comparison by calculating the e-values reported by Propositions 5.3 and 5.6.
Example 5.7.
Consider a two-round betting with borrowings $\beta_{1}=\beta_{2}=1$ and bets $\lambda_{1}=\lambda_{2}=1/2$. A sub-liabilities $\tilde{L}_{2}=\beta_{1}+(2-X_{1})\beta_{2}$ is defined to penalize borrowing when already at a loss. We can see from Table 1 that it is possible that the e-value corresponding to the sub-net wealth is larger.
$X_{1}$ $X_{2}$ $W_{2}$ $N_{2}$ $\frac{N_{2}+1}{2}$ (e-value) $\tilde{N}_{2}$ $\frac{\tilde{N}_{2}+3}{4}$ (e-value)
-1 -1 1 -1 0 -3 0
-1 1 3 1 1 1 1
1 -1 2 0 0.5 0 0.75
1 1 6 4 2.5 4 1.75
Table 1: Net wealth, sub-net wealth and their e-values.
6 Borrowing as Averaging
In this section, we demonstrate a case where the stopped e-value that underlies Proposition 5.3, $(N_{\tau}-N_{\min})/(1-N_{\min})$, is equivalent to an average over several e-values, each of which can be understood without the borrowing scenario we introduced.
We assume that Statistician is committed to the following “bet and save” strategy: they periodically set aside their current gross wealth into savings and only put their newly borrowed money on the table. Formally, we consider a sequence of strictly increasing bounded stopping times $\{\tau_{n}\}$ with $\tau_{0}=0$, which we call “save times”. We further assume that the borrowings process $\{\beta_{t}\}$ satisfies, for all $t\geqslant 1$, $\beta_{t}\geqslant 0$ and
$\bigcap_{n=0}^{\infty}\{\tau_{n}+1\neq t\}\subseteq\{\beta_{t}=0\},$ (15)
meaning that Statistician only borrows before the bets on $X_{\tau_{n}+1}$, $n=0,1,\dots$, and between two consecutive borrows, their wealth evolves from $W_{\tau_{n}}+\beta_{\tau_{n}+1}$ to $W_{\tau_{n+1}}$.
Let us further assume, crucially, that the bets $\lambda_{\tau_{n}+1},\dots,\lambda_{\tau_{n+1}}$, i.e., those that occur during the aforementioned period between two consecutive borrows, are chosen such that the wealth $W_{\tau_{n}}$ is never risked; that is, $W_{t}(\omega)\geqslant W_{\tau_{n}(\omega)}(\omega)$ for any $\omega\in\Omega$ and $t\in[\tau_{n}(\omega),\tau_{n+1}(\omega)]$. This effectively amounts to the wealth $W_{\tau_{n}}$ accrued being set aside from upcoming rounds of bets.
Then, Statistician can then compute an e-value for each time period $\tau_{n}$ to $\tau_{n+1}$: in each of such period, the proceeds $W_{\tau_{n+1}}-W_{\tau_{n}}$ are generated solely by the $\beta_{\tau_{n}+1}$ that Statistician initially borrowed. Therefore, an e-value can be defined by the ratio between the proceeds $W_{\tau_{n+1}}-W_{\tau_{n}}$ and the initial capital $\beta_{\tau_{n}+1}$,
$E_{n}=\frac{W_{\tau_{n+1}}-W_{\tau_{n}}}{\beta_{\tau_{n}+1}}.$ (16)
Indeed, a direct calculation shows that $\mathbb{E}(W_{\tau_{n+1}}-W_{\tau_{n}}|\mathcal{F}_{\tau_{n}})=\beta_{\tau_{n}+1}$. Note that the e-values $E_{n}$ can be understood without reasoning about borrowing, as no borrowing occurs besides $\beta_{\tau_{n}+1}$, which is treated here as an initial capital.
In order to connect to Proposition 5.3, we further assume that Statistician stops such procedure at a final stopping time $\tau_{B}$ when reaching a non-random total liabilities
$L=L_{\tau_{B}}=\sum_{n=0}^{B-1}\beta_{\tau_{n}+1}.$ (17)
The $B$ here, the total number of such borrow-bet-save periods, is possibly random (technically, a finite stopping time on the filtration $\{\mathcal{F}_{\tau_{n}}\}_{n\geqslant 0}$). Statistician so far has obtained $B$ e-values,
$E_{0},\dots,E_{B-1},$ (18)
each $E_{n}$ satisfying $\mathbb{E}_{1/2}(E_{n}|\mathcal{F}_{\tau_{n}})=1$. Note that each $\beta_{\tau_{n}+1}$ is $\mathcal{F}_{\tau_{n}}$-measurable, a weighted average over (18) can be taken, where each $E_{n}$ receiving weight $\beta_{\tau_{n}+1}$, to obtain another e-value,
$E=\frac{\sum_{n=0}^{B-1}\beta_{\tau_{n}+1}E_{n}}{\sum_{n=0}^{B-1}\beta_{\tau_{n}+1}}.$ (19)
To see that $E$ is indeed an e-value,
$\mathbb{E}_{1/2}(E|\mathcal{F}_{\tau_{B}})=\frac{\mathbb{E}_{1/2}\{\sum_{n=0}^{B-1}\mathbb{E}_{1/2}(\beta_{\tau_{n}+1}E_{n}|\mathcal{F}_{\tau_{n}})|\mathcal{F}_{\tau_{B}}\}}{L}=\frac{\sum_{n=0}^{B-1}\mathbb{E}_{1/2}(\beta_{\tau_{n}+1}|\mathcal{F}_{\tau_{B}})}{L}=1.$ (20)
And we can simplify the expression for $E$ by
$E=\frac{\sum_{n=0}^{B-1}\beta_{\tau_{n}+1}\frac{W_{\tau_{n+1}}-W_{\tau_{n}}}{\beta_{\tau_{n}+1}}}{L}=\frac{W_{\tau_{B}}-1}{L}.$ (21)
Alternatively, let us compute the minimum net wealth that Statistician can ever have over the course until $\tau_{B}$. Since the most they can lose from $\tau_{n}$ to $\tau_{n+1}$ is the amount $\beta_{\tau_{n}+1}$ borrowed, the following is an almost sure lower bound of Statistician’s net wealth:
$N_{\min}=\operatorname*{ess\,inf}\left(1-\sum_{n=0}^{B-1}\beta_{\tau_{n}+1}\right)=\operatorname*{ess\,inf}(1-L)=1-L.$ (22)
Therefore, the e-value that underlies Proposition 5.3 is
$E^{\prime}=\frac{N_{\tau_{B}}-N_{\min}}{1-N_{\min}}=\frac{M_{\tau_{B}}-L-(1-L)}{L}=\frac{W_{\tau_{B}}-1}{L}.$ (23)
We thus conclude that $E=E^{\prime}$, i.e., the two different perspectives yield the same e-value.
7 The Futile Borrower
We have seen so far that while borrowing does lead to evidence via gambling wealth or net wealth, the evidence is to be discounted by how much is borrowed, either factually or counterfactual. We thus ask a higher-level question: does Statistician obtain any better evidence by borrowing? Since evidence is defined in Section 3 with two extra discount factors $a$ and $b$, it is necessary to standardize all tail evidences to, say, $(1,0)$-tail evidences to which e-values belong for a fair comparison.
We study the general case in this section that encompasses the coin betting set-up discussed so far. For a random variable $Y$ and a probability measure $\mathbb{P}$ on the measurable space $(\Omega,\mathcal{A})$, we write the expected value of $Y$ under $\mathbb{P}$ as $\mathbb{E}_{\mathbb{P}}Y$.
Consider a general random variable $Y$ on $(\Omega,\mathcal{A})$, understood as the outcome of a bet when risking unit initial capital. Then, the (net) leveraged outcome of the bet, when borrowing $\beta$ and putting $1+\beta$ on the same bet, is the random variable $(1+\beta)Y-\beta$. We consider a set of random variables on $(\Omega,\mathcal{A})$ that is closed under the leverage map: if $Y\in\mathcal{V}$ and $\beta\neq 1$, then $(1+\beta)Y-\beta\in\mathcal{V}$, and define the following concept.
Definition 7.1 (Leverage invariance).
A functional $f:\mathcal{V}\to{\mathbb{R}}$ is said to be leverage invariant over $\mathcal{V}$ if
$f((1+\beta)Y-\beta)=f(Y)$ (24)
for all $Y\in\mathcal{V}$ and $\beta\neq-1$.
For example, in the modern portfolio theory, the Sharpe ratio that measures expected return relative to risk is a classical example of a leverage invariant quantity. To wit, let $\mathbb{P}$ be any probability measure on $(\Omega,\mathcal{A})$ and $L^{2}(\mathbb{P})$ be the set of all (non-constant) square-integrable random variables. It is easy to see that the functional
$\operatorname{Sharpe}(Y)=\frac{\mathbb{E}_{\mathbb{P}}Y-1}{\sqrt{\mathbb{E}_{\mathbb{P}}((Y-\mathbb{E}_{\mathbb{P}}Y)^{2})}}$ (25)
is leverage invariant over $L^{2}(\mathbb{P})$. See e.g. Brugière [2020, Proposition 6.1.1] for more on this classical concept.
In our case, let us consider two probability measures, $\mathbb{P}$ and $\mathbb{Q}$, that represent the null and alternative distributions. On the set of all $\mathbb{Q}$-integrable random variables $L^{1}(\mathbb{Q})$, we define the functional
$\mathcal{E}(Y)=\sup_{a,b}\left\{\mathbb{E}_{\mathbb{Q}}(aY+b):\mathbb{E}_{\mathbb{P}}(aY+b\geqslant x)\leqslant x^{-1}\right\},$ (26)
that is, if we standardize $Y$ by an affine transformation such that it becomes a $(0,1)$-tail evidence against $\mathbb{P}$, $\mathcal{E}(Y)$ is the maximum expected value under $\mathbb{Q}$. We have the following statement.
Proposition 7.2.
The functional $\mathcal{E}$ is leverage invariant over $L^{1}(\mathbb{Q})$.
Proof.
Note that
$\mathcal{E}(Y)=\sup_{a,b}\left\{a\mathbb{E}_{\mathbb{Q}}(Y)+b:\mathbb{E}_{\mathbb{P}}(aY+b\geqslant x)\leqslant x^{-1}\right\}.$ (27)
And
$\mathcal{E}((1+\beta)Y-\beta)=\sup_{a,b}\left\{a(1+\beta)\mathbb{E}_{\mathbb{Q}}(Y)+b-a\beta:\mathbb{E}_{\mathbb{P}}(a(1+\beta)Y+b-a\beta\geqslant x)\leqslant x^{-1}\right\}.$ (28)
Since $(a,b)\mapsto(a(1+\beta),b-a\beta)$ is an invertible linear transformation if $\beta\neq-1$, the two suprema above are essentially the same supremum merely under different parameterizations. Thus we conclude that $\mathcal{E}(Y)=\mathcal{E}((1+\beta)Y-\beta)$. ∎
We remark that the same result holds true if we replace $\mathbb{E}_{\mathbb{Q}}(Y)$ by any linear score function on random variables. The leverage invariance of $\mathcal{E}$ implies that, in one-round testing by betting games, it is fundamentally impossible to obtain a “better” evidence by only borrowing (without changing the betting structure) and announcing the resulting net wealth. This happens, we note, only under the criterion that the “goodness” of an evidence is evaluated by its expected value after a proper rescaling to a (“standardized”) $(0,1)$-tail evidence. To put it another way, no matter how one devises a borrowing scheme, the resulting net wealth, if standardized, has the same expected value under any fixed alternative distribution. Of course, the expected value of standardized evidence is in no way the only evaluation criterion for the quality of an evidence. Non-linear utility functions can be applied, for example: many authors such as Kelly, Breiman and Shafer [2021] argue that the expected log wealth is more meaningful in the borrow-free setting. This, however, fails to be compatible with the possible non-positivity of net wealth we study in this paper.
8 Almost Supermartingales and Bargained Bets
Still working with a general scenario, we equip the measurable space $(\Omega,\mathcal{A})$ with a fixed probability $\mathbb{P}$ and an arbitrary filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$ in this section.
Robbins and Siegmund [1971] define the following class of stochastic processes.
Definition 8.1 (Almost supermartingale).
A process $\{Z_{t}\}_{t\geqslant 0}$ adapted to $\{\mathcal{F}_{t}\}_{t\geqslant 0}$ is said to be an almost supermartingale with respect to predictable nonnegative processes $\{b_{t}\}_{t\geqslant 1},\{\xi_{t}\}_{t\geqslant 1}$ and $\{\zeta_{t}\}_{t\geqslant 1}$, if
$\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})=Z_{t-1}(1+b_{t})+\xi_{t}-\zeta_{t}$ (29)
for all $t\geqslant 1$.
One can see that any integrable stochastic process $\{Z_{t}\}$ adapted to $\{\mathcal{F}_{t}\}$ is an almost supermartingale with respect to some appropriately chosen triplet of predictable processes. For example, $b_{t}=0$, $\xi_{t}=(\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})-Z_{t-1})^{+}$, and $\zeta_{t}=(\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})-Z_{t-1})^{-}$. The following maximal inequality holds for nonnegative almost supermartingales.
Proposition 8.2 (Proposotion 2 of Robbins and Siegmund [1971]).
Suppose $\{Z_{t}\}_{0\leqslant 1\leqslant T}$ is nonnegative and satisfies Definition 8.1. Let $Z^{*}_{T}=\max_{t\leqslant T}Z_{t}$. Then,
$\mathbb{P}(Z_{T}^{*}\geqslant x)\leqslant x^{-1}\left(\mathbb{E}Z_{0}+\sum_{t=1}^{T}\mathbb{E}\xi_{t}\right)+\sum_{t=1}^{T}\mathbb{E}b_{t}.$ (30)
Clearly, the inequality holds for $T=\infty$ as well. The inequality can be reduced to various maximal inequalities in the literature, for example Ville’s inequality for nonnegative supermartingales (taking $b_{t}=\xi_{t}=0$) and Doob’s inequality for nonnegative submartingales (taking $b_{t}=\zeta_{t}=0$). To see, for example, that our Proposition 4.4 also follows from Proposition 8.2, simply note that the wealth process $\{W_{t}\}$ is a nonnegative almost supermartingale with respect to $b_{t}=0$, $\xi_{t}=\beta_{t}^{+}$, and $\zeta_{t}=\beta_{t}^{-}$.
All these examples, however, are special cases of Proposition 8.2 with the $\{b_{t}\}$ terms set to zero. When could the $\{b_{t}\}$ be non-zero? Let us consider the scenario where Casino misprices the bet – either by mistake or on purpose, for example, when Statistician has successfully bargained for a higher return per dollar at stake.
In all our previous exposition, the wealth of Statistician evolves as
$W_{t}=(W_{t-1}+\beta_{t})\cdot(1+\lambda_{t}X_{t}),$ (31)
because Casino offers double-or-nothing bets, on both heads and tails, in which case the bets are fairly priced to the effect that no arbitrage opportunities exist for Statistician to win risk-free money. Let us instead consider the bets offered on $X_{t}$ being $(2+2b_{t})$-times-or-nothing for some $b_{t}>0$: for any dollar placed on either heads or tails333 Of note, these mispriced bets we consider here are offered two-sided in symmetry. If only one side is mispriced, or if two sides are priced differently (which normally would fairly price a non-$\mathbb{P}_{1/2}$ null), the condition for the existence of arbitrage shall be different accordingly., Casino will pay back $2+2b_{t}$ dollars if correctly guessed. In this case, we note, Statistician is able to arbitrage by putting $1/2$ on heads and $1/2$ on tails, increasing wealth from 1 to $1+b_{t}$ regardless of the actual distribution $\mathbb{P}_{p}$ and the actual outcome of the coin toss $X_{t}$.
Statistician’s wealth in this mispriced case evolves as
$W_{t}=(W_{t-1}+\beta_{t})\cdot(1+b_{t})(1+\lambda_{t}X_{t}),$ (32)
where $\lambda_{t}\in[-1,1]$ as before. The multiplier $(1+b_{t})(1+\lambda_{t}X_{t})$ is interpreted as follows: Statistician allocates $|\lambda_{t}|$ portion of their wealth into risky bets on either heads ($\lambda_{t}>0$) or tails ($\lambda_{t}<0$), paying off $(1+b_{t})|\lambda_{t}|+(1+b_{t})\lambda_{t}X_{t}$ as per Casino’s protocol; and $1-|\lambda_{t}|$ of their wealth on the risk-free 50-50 combination of bets on heads and tails, paying off $(1-|\lambda_{t}|)(1+b_{t})$, as it does not make sense for Statistician to keep this idle cash. In total, therefore, the wealth is to be multiplied by $(1+b_{t})(1+\lambda_{t}X_{t})$.
It is easy to see that the wealth process $\{W_{t}\}$ is a nonnegative almost supermartingale with respect to $b_{t}$, $\xi_{t}=(1+b_{t})\beta^{+}_{t}$, and $\zeta_{t}=(1+b_{t})\beta^{-}_{t}$. To turn it via Proposition 8.2 into an evidence, let us slightly extend our previously defined concept of $(a,b)$-sequential tail evidence (Definition 3.3).
Definition 8.3 (Approximate sequential tail evidence).
If a process $\{E_{t}\}$ adapted to $\{\mathcal{F}_{t}\}$ satisfies, for all $x>b$,
$\mathbb{P}_{1/2}\left(\sup_{t}E_{t}\geqslant x\right)\leqslant\frac{a}{x-b}+c,$ (33)
we say that it is an $(a,b,c)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
Then, we immediately have the following.
Proposition 8.4.
If
$B=\sup_{t}\mathbb{E}_{1/2}\sum_{i=1}^{t}(1+b_{i})\beta_{i}^{+}<\infty,\quad C=\sum_{t=1}^{\infty}b_{t}<\infty,$ (34)
the wealth process $\{W_{t}\}$ is a $(1+B,0,C)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
As can be already seen from the new definition of $(a,b,c)$-sequential tail evidence in the statement, the evidence that the wealth process $\{W_{t}\}$ generates in this case is of lower quality; there is always a $C$ probability no matter how large the gross wealth has been accrued. That is, as long as $C\geqslant 0.05$, the corresponding $0.05$-level sequential test is trivial. We thus seek to develop evidence with net wealth instead, as we have done in Section 5. However, if we still define the net wealth as $W_{t}-(\beta_{1}+\dots+\beta_{t})$, immediately we run into trouble: since arbitrage opportunities exist, Statistician can borrow an unbounded amount of money, put it all in the risk-free 50-50 combination of bets on heads and tails, and take a net gain of $b$ multiplied by the borrowed amount. Not only the wealth but the net wealth increases unbounded and risk-free, with absolutely no insight gained into the verity of $\mathbb{P}_{1/2}$ as the growth is completely independent of the outcomes. The crux of the matter lies in our earlier assumption that Statistician can borrow without interest. When a risk-free instrument exists, financial literature tells us that the interest of borrowing shall be set the same as the risk-free return rate to prevent pathological possibilities in the system.
We thus define the compound interest liabilities process,
$L_{0}=0,\quad L_{t}=(1+b_{t})(L_{t-1}+\beta_{t}),$ (35)
and the net wealth as $N_{t}=W_{t}-L_{t}$.
Proposition 8.5.
The equation
$\underbrace{\frac{W_{t}}{(1+b_{1})\dots(1+b_{t})}}_{W_{t}^{\prime}}=\underbrace{\frac{N_{t}}{(1+b_{1})\dots(1+b_{t})}}_{N_{t}^{\prime}}+\underbrace{\frac{L_{t}}{(1+b_{1})\dots(1+b_{t})}}_{L_{t}^{\prime}}$ (36)
is the Doob decomposition of $\{W_{t}^{\prime}\}$ into the martingale $\{N_{t}^{\prime}\}$ and the predictable process $\{L_{t}^{\prime}\}$.
Proof.
It suffices to show that $L_{t}^{\prime}-L_{t-1}^{\prime}=\mathbb{E}_{1/2}(W_{t}^{\prime}|\mathcal{F}_{t-1})-W_{t-1}^{\prime}$. Note that
$L_{t}^{\prime}-L_{t-1}^{\prime}=\frac{(1+b_{t})(L_{t-1}+\beta_{t})-(1+b_{t})L_{t-1}}{(1+b_{1})\dots(1+b_{t})}=\frac{\beta_{t}(1+b_{t})}{(1+b_{1})\dots(1+b_{t})}$ (37)
$=$ $\frac{(W_{t-1}+\beta_{t})(1+b_{t})-(1+b_{t})W_{t-1}}{(1+b_{1})\dots(1+b_{t})}=\mathbb{E}_{1/2}(W_{t}^{\prime}|\mathcal{F}_{t-1})-W_{t-1}^{\prime},$ (38)
concluding the proof. ∎
Therefore, if one places a lower bound on the “adjusted” net wealth process $\{N_{t}^{\prime}\}$, which is a martingale, all results in Section 5 can be straightforwardly replicated.
9 Summary
We study in this paper the extension of “testing by betting”, a central topic in game-theoretic statistics, into the setting where borrowing is allowed. Our conclusion is twofold. There indeed exist multiple formulations of evidence (“tail evidence”, as defined by rare events under the model to be rejected) via either the total or the net wealth after betting, under a range of assumptions that limit the borrowed volume. On the other hand, however, we see that if net wealth after betting is ever to be counted as evidence, it is subject to a fundamental invariance principle with respect to borrowing. This, we note, shall not be seen as a knock-down refutation against borrowing in testing-by-betting, since at least it is possible to produce larger “standardized evidence” (e-values) for some possible outcomes. We separately discuss the effect of mispricing bets (bargaining) and the correct ways to quantify evidence correspondingly, which is an issue of independent interest in testing-by-betting completely formulable without introducing borrowing, but it nonetheless makes an interesting variant in our current paper since infinite arbitrage can be made with borrowing.
Our study also leaves numerous issues for future work. First, we often define a sequential evidence by penalizing an upper bound (either almost surely or in expectation) of borrowing, as opposed to the current amount borrowed. Is it possible to define a sequential evidence that depends on borrowing only via the current liabilities? Second, while we do construct a toy example (Table 1) where borrowing produces a larger e-value, it fails to be very insightful and it remains unanswered how in general borrowing can be beneficial evidence-wise (we do know, in contrast, that borrowing can be beneficial financially when one is very sure about the underlying truth, but evidence asks for an arguably stronger condition). Finally, the principle of “leverage invariance” for borrowing is proven to upper bound the expected value of a standardized evidence under an alternative distribution. It remains unclear if there are better arguments to be made in this respect. Indeed, in the borrowing-free case one usually considers the expected logarithm of an e-value under an alternative distribution, which fails to be compatible with borrowing due to possible negativity. Replacing the logarithm with some other concave utility functions that allow negative values, we believe, could be the next thing to consider.